OpenStudy (anonymous):

My answer is c Solve the combined inequality and select the answer that best describes the graph of the solution. 3x - 7 < -10 or 5x + 2 > 22 Open dot at -1 shaded to the left, open dot at 4 shaded to the right Closed dot at -1 shaded to the left, closed dot at 5 shaded to the right Open dot at 1 shaded to the left, open dot at 5 shaded to the right None of the above

OpenStudy (anonymous):

i will bet you can do this

OpenStudy (anonymous):

hahahahaha is that a joke

OpenStudy (anonymous):

it would be a closed dot.

OpenStudy (anonymous):

for example if \(3x-7<-10\) the add then 7 to both sides and get \[3x<-3\] divide by 3 and get \[x<-1\]

OpenStudy (anonymous):

what the hell just happened??

OpenStudy (anonymous):

all "dots" are open for < and for >

OpenStudy (anonymous):

omg you said the H word hahahahahahahhahaahhaahah

OpenStudy (anonymous):

only "closed" for \(\leq \)

OpenStudy (anonymous):

\[3x-7 < -10 \rightarrow 3x < -3 \rightarrow x < -1\]\[5x+2>22 \rightarrow 5x > 20 \rightarrow x >4 \]If the function only exists for x's SMALLER than -1 and x's BIGGER than 4, then what kind of dot would represent that? @InsanelyChaotic

OpenStudy (anonymous):

#confused

OpenStudy (anonymous):

where is my latex?????

OpenStudy (anonymous):

satan ate it @satellite73

OpenStudy (anonymous):

omg latex as in cond***

OpenStudy (anonymous):

no no latex as in the code i use to write \[\leq\]

OpenStudy (anonymous):

oh;)

OpenStudy (anonymous):

closed dots used for \(\leq \) and for \(\geq\)

OpenStudy (anonymous):

okay!!:)

OpenStudy (anonymous):

open dots for \(<\) and \(>\)

OpenStudy (anonymous):

gotcha

OpenStudy (anonymous):

that is easy enough to remember, not much thinking required

OpenStudy (anonymous):

somehow the latex disappeared now it is back

OpenStudy (anonymous):

so i can enjoy safe math symbols

OpenStudy (anonymous):

still confused on the result. i cant understand without pics

OpenStudy (anonymous):

hahahahah oh my god that's funny

OpenStudy (anonymous):

you have to solve both inequalities, then put the word "or" between them the first one gives \(x<-1\) the second one gives \(x>4\)

OpenStudy (anonymous):



OpenStudy (anonymous):

so open dot at -1, shaded left, open dot at 4, shaded right who said art was dead?

OpenStudy (anonymous):

PERFECT PERFECT PERF PERF

OpenStudy (anonymous):

Open dot means that the "<" or ">" signify a non-inclusive interval. Observe: @InsanelyChaotic